OpenStudy (anonymous):

10 cm^3 = ___ m^3 [Steps drawn please! D: ]

OpenStudy (anonymous):

= (0.1m)^3 = (0.1^3)m^3

OpenStudy (anonymous):

=0.001m^3

OpenStudy (anonymous):

=1/(10^5)

OpenStudy (anonymous):